# VecCity: A Taxonomy-guided Library for Map Entity Representation Learning [Experiment, Analysis & Benchmark]

### v1

@@ 1. Introduction
In the era of mobile internet, electronic maps have become a foundational platform for a wide range of applications, including intelligent transportation systems and location-based services. An electronic map consists of map entities, such as points of interest (POIs), road segments, and land parcels. These entities encapsulate complex geospatial data, spatial relationships, and geometric-topological structures, presenting significant challenges for data representation. The effective representation of map entities has become a critical and enduring research focus within spatiotemporal data analysis and geographic information systems (GIS).
In the literature, early studies represented map entities as records in relational databases or as files containing geographic information attributes, such as PostGIS (Team, 2023), Shapefiles (Environmental Systems Research Institute, 1998), and GeoJSON (Butler et al., 2016). With this representation approach, application services relying on map entities are required to develop task-specific models or algorithms to process the entities’ attributes and extract geographic relationships within or between them. However, the task-specific nature of these data processing models restricts their generalizability and reusability across different applications and scenarios.
Figure 1. ArXiv submissions on MapRL over time.
In recent years, pre-trained representation learning has emerged as a powerful approach for generating versatile, task-independent data representations. It has achieved remarkable success across various domains, including natural language processing (NLP) (Radford et al., 2018; Mikolov et al., 2013; Peters et al., 2018; Pennington et al., 2014) and computer vision (CV) (Chen et al., 2020; He et al., 2020, 2016; Dosovitskiy et al., 2021; Vincent et al., 2010). In the field of spatiotemporal data analysis, a growing number of studies have embraced pre-trained representation learning methods to construct generic representations of map entities (Mai et al., 2022; Chen et al., 2024a; Zhang et al., 2024; Li et al., 2024b; Ning et al., 2024). These efforts have given rise to an emerging research area known as Map Entity Representation Learning (MapRL). As shown in Fig. 1, the number of MapRL papers has been steadily increasing, reflecting the growing interest and potential impact of this field. However, despite these rapid advancements, several structural challenges require further exploration to unlock the potential of MapRL.
@@
Challenge 2: Lack of Standardized Benchmark. Although numerous MapRL models have been proposed in the literature, their performance is often evaluated on different datasets and under varying experimental settings. Unlike the CV and NLP fields, this domain lacks standardized datasets and benchmarks. This absence hinders fair performance comparisons between models and limits the ability of researchers to establish guiding design principles through consistent and rigorous model evaluation.
To address the first challenge, we propose a novel taxonomy for MapRL models, namely method-based taxonomy, which organizes the essential components of an MapRL model into four key elements: Map Data, Encoder Models, Pre-training Tasks, and Downstream Tasks. The map data is further divided into Map Entity Data and Auxiliary Data. Encoder models are categorized into three types: Token-based, Graph-based, and Sequence-based models, each supporting different pre-training tasks. A key insight of our taxonomy is that the two core elements—Encoder Models and Pre-training Tasks—are not tightly coupled with specific map entity types. Only the downstream tasks depend on the type of map entity. The method-based taxonomy represents a significant departure from the traditional classification of MapRL models, i.e., type-based taxonomy, which primarily focuses on map entity types.
To address the second challenge, we propose an MapRL model development toolkit based on our taxonomy, namely the VecCity library. The library organizes map entities and auxiliary data (i.e., the map data element) into three types of atomic files. Using these atomic files, we format data from nine cities into a unified structure. the VecCity library further provides four interface functions to standardize the implementation processes (i.e., the encoder model, pre-training task, and downstream task elements) of MapRL models. Leveraging these interfaces, we have implemented 21 mainstream MapRL models. Additionally, we conduct a comprehensive evaluation of these models across the datasets from the nine cities, offering valuable insights to guide future research.
@@
  • Novel Taxonomy. We present a novel taxonomy of MapRL models that goes beyond the traditional classification based on map entities, offering a unified classification system applicable to MapRL models across various types of map entities.
  • Easy-to-use Toolkit Library. Guided by our taxonomy, we propose a toolkit, VecCity, which unifies the development process of MapRL models. To the best of our knowledge, VecCity is the first unified model toolkit specifically designed for MapRL.
  • Standard Benchmark. Using VecCity, we implement 21 mainstream MapRL models and evaluate their performance on datasets from nine cities. This work establishes the first standard benchmark for all types of MapRL models and provides valuable insights to guide future research.
Figure 2. The framework of a MapRL model.
@@ 2. Taxonomy of MapRL
In this section, we present the proposed method-based taxonomy. Figure 2 outlines the general framework for MapRL. The training process of a typical MapRL model consists of two key stages: pre-training and fine-tuning. During pre-training, the MapRL model takes map data as input, using an encoder model to transform them into representation vectors. The encoder’s parameters are optimized through pre-training tasks. In the fine-tuning stage, downstream tasks are employed to further refine the encoder. This pipeline defines four core components of an MapRL model: Map Data, Encoder Models, Pre-training Tasks, and Downstream Tasks.
@@ 2.1.1. Map Entities
As illustrated in Fig. 3, an electronic map consists of three geometric elements: Points, Polylines, and Polygons, each representing a distinct type of map entity.
  • Points / POI. A point element represents a Point of Interest (POI), identified by a location coordinate, i.e., $(latitude,longitude)$.
  • Polylines / Road Segment. A polyline element represents a Road Segment, composed of a sequence of connected line segments, with each endpoint defined by location coordinates.
  • Polygons / Land Parcel. A polygon element represents a Land Parcel, defined by a set of connected line segments forming a closed boundary.
For each map entity, a set of features describes its properties. For instance, the category feature of a POI differentiates the services it offers, such as scenic spots, transit stations, or shopping facilities. Road segments are characterized by features like speed limits, lane counts, road grade, etc. Based on these definitions, we formally define electronic maps and map entities as follows.
@@ 2.1.2. Auxiliary Data
Trajectory data can be divided into two classes based on the type of location samples $l_{tr_{k}}$. $i$) Check-in Trajectories: These use POIs as samples and are collected through social networks or location-based service (LBS) apps. They log a user’s location only upon check-in, resulting in a low sampling rate, often requiring days to capture a single point. MapRL models leverage these trajectories to infer users’ personal preferences. $ii$) Coordinate Trajectories: These rely on latitude and longitude coordinates as samples. Collected by intelligent transportation systems (ITS) through GPS devices or mobile phones, they provide high-frequency sampling along with real-time speed and direction data. These trajectories are typically aligned with road segments using map matching algorithms (Yang and Gidofalvi, 2018) and facilitate route reconstruction and dynamic state estimation for road segments.
@@
A relation network represents map entities as a directed graph $G=\{V,\bm{E}\}$, where $V=\{v_{1},\ldots,v_{I}\}$ is the set of vertices corresponding to map entities. $\bm{E}\in\mathbb{R}^{I\times I}$ is an adjacency matrix, with each element $e_{ij}$ indicating the relationship from $v_{i}$ to $v_{j}$.
Network relations can be categorized into two types: $i$) Geographical Relations (GR): These leverage spatial relationships between map entities to construct the adjacency matrix. Examples include connections between road segments in a road network (Zhang and Long, 2023) and distances between POIs (Zhao et al., 2017; Du et al., 2019). $ii$) Social Relations (SR): These capture user behaviors, such as mobility trajectories, to form adjacency matrices. A common example is Origin-Destination (OD) flows, widely used in MapRL models (Zhang et al., 2020; Zhou et al., 2023), where the number of trajectories originating from $v_{i}$ and terminating at $v_{j}$ defines the value of $e_{ij}$ in the adjacency matrix.
@@ 2.2. Encoder Models and Pre-training Tasks
Encoder models and pre-training tasks are two core components of a pre-trained MapRL model, responsible for converting input map data into representation vectors for map entities. Our taxonomy classifies encoder models into three types: token-based, graph-based, and sequence-based, designed to model features, relation networks, and trajectory sequences of map entities, respectively. Each encoder type corresponds to different pre-training tasks tailored to its structure and purpose.
@@ 2.2.1. Token-based Models
Encoders. The token-based encoder generates representation vectors directly from the features of map entities. For discrete features, it applies one-hot encoding, and for continuous features, the value range is divided into consecutive bins, each represented by a one-hot code. Given $K$ features for a map entity, let the one-hot code of the $k$-th feature be $\bm{f}_{k}\in\{0,1\}^{F_{k}}$ (where $F_{k}$ is the dimensionality). The encoder employs a learnable embedding matrix $\bm{R}_{k}\in\mathbb{R}^{F_{k}\times d}$ for each feature to generate the entity’s representation:
(1) $\small\bm{r}=\left\|{}_{k=1}^{K}\bm{f}_{k}^{\top}\cdot\bm{R}_{k}\right.,$
where $\|_{k=1}^{K}$ denotes the concatenation of all features’ embeddings.
In the literature, the novelty of existing token-based methods lies in how effectively they extract valuable features from map data. These features can be categorized into three types. $i$) Spatial Features: These capture the locations and spatial structures of map entities. Spatial coordinates are the most fundamental spatial features. Hier (Shimizu et al., 2020) incorporates spatial structures by embedding hierarchical spatial grids into the representations of rasterized maps. $ii$) Temporal Features: A map entity’s characteristics can vary at different times of the day. For example, a restaurant exhibits different behaviors during working hours versus lunchtime. To capture these variations, many token-based encoders divide the day into segments (e.g., hourly or half-hourly) and assign each segment a unique embedding to represent temporal features (Wan et al., 2021; Zhao et al., 2017; Gong et al., 2023; Ma et al., 2024; Liu et al., 2016). $iii$) Semantic Features: These reflect the intrinsic properties of map entities, such as the functional types of POIs or the road types of segments. Some land parcel MapRL methods further enhance semantic features by counting the number of POI categories within a parcel (Zhang et al., 2022).
@@
  • Token Relation Inference (TokRI). This task aims to predict the relationship between two map entities based on their representation vectors. Specifically, given the representations $\bm{r}_{i}$ and $\bm{r}_{j}$ of two entities, the TokRI task predicts their relationship as $\hat{y}_{ij}=\phi(\bm{r}_{i},\bm{r}_{j})$, where $\phi(\cdot)$ is the prediction function, and $\hat{y}_{ij}\in[0,1]$ is the predicted probability of a relation existing. The encoder parameters are optimized during pre-training using cross-entropy loss.
@@
  • Token Relation Contrastive Learning (TRCL). This task ensures that entities with closer relationships are more likely to have similar representations. Given a representation sample $\bm{c}_{i}$, another sample $\bm{c}_{j}$ is treated as a positive sample if the corresponding entities $v_{i}$ and $v_{j}$ share a close relation; otherwise, $\bm{c}_{j}$ is treated as a negative sample. The encoder parameters are optimized using a contrastive learning loss, such as the InfoNCE loss (Oord et al., 2018).
@@
  • Augmented Token Contrastive Learning (AToCL). These tasks use data augmentation techniques to generate different views of a given entity’s representation. A contrastive learning loss is then applied to maximize the similarity between views of the same entity while minimizing the similarity between different samples.
@@
For TokRI and TRCL tasks, the novelty of existing models stems from their methods to constructing token relations. The novelty of existing TokRI and TRCL tasks lies in their approaches to constructing token relations, which can be categorized into two methods: $i$) Geographical Distance-Based Relations: This method uses spatial proximity to establish token relations. In TokRI tasks, RN2Vec (Wang et al., 2020) defines a relationship between entities if their geographical distance is below a preset threshold. In TRCL tasks, Teaser (Zhao et al., 2017) selects the $K$ nearest entities as positive samples for contrastive learning. $ii$) Auxiliary Data-Based Relations: This method leverages additional data, such as trajectories, to define relations. In TokRI tasks, P2Vec (Feng et al., 2017) and Tale (Wan et al., 2021) treat a trajectory as a sentence and the entities within it as words, using the Continuous Bag of Words (CBOW) model (Mikolov et al., 2013) to learn representations for “words” (map entities). In TRCL tasks, Teaser (Zhao et al., 2017) employs the SkipGram model (Mikolov et al., 2013) with negative sampling to capture contextual relations along trajectories. When trajectory data is unavailable, “virtual trajectories” can be generated via random walks on relation networks (Wang and Li, 2017). Beyond trajectories, ZEMob (Yao et al., 2018) and DeepMove (Zhou and Huang, 2018) leverage OD flows between entities to identify relationships.
@@
For AToCL tasks, the novelty of existing models lies in their methods for generating augmented samples. For example, ReMVC (Zhang et al., 2022) and ReDCL (Li et al., 2023) augment data by replacing or dropping parts of the anchor entity’s features.
@@ 2.2.2. Graph-based Models
Encoders. Graph-based encoders leverage auxiliary data in the form of relation networks (see Definition 3) to capture relationships among map entities. Graph-based encoders typically use the representations generated by token-based encoders as inputs and apply a graph neural network (GNN) to produce refined representation vectors. Specifically, given the map entity set $V=\{v_{1},\ldots,v_{I}\}$, the embeddings generated by the token-based encoder are denoted as $\bm{R}=\{\bm{c}_{1},\ldots,\bm{c}_{I}\}$. The graph-based encoder then takes $\bm{R}$ along with the relation network $G$ to generate the final representations:
(2) $\bm{H}=\left\{\bm{h}_{1},\ldots,\bm{h}_{i},\ldots,\bm{h}_{I}\right\}=\mathrm{GNN}\left(\bm{R},G\right),$
where $\bm{h}_{i}$ is the representation vector for the $i$-th map entity, and $\mathrm{GNN}(\cdot)$ is the GNN-based encoder model. Compared to token-based models, graph-based encoders incorporate the relation networks into the entity representations, enriching their expressiveness.
@@
Existing research on graph-based encoders focuses on developing GNNs capable of capturing complex relationships among map entities. Early studies applied standard GNN models, such as graph convolutional networks (Yu et al., 2018; Wu et al., 2019) and graph attention networks (Song et al., 2022; Zhang et al., 2019b), to explore graphical (Chang et al., 2023; Jiang et al., 2023; Jiao et al., 2020) and social relations (Zhang et al., 2020; Liu et al., 2020). Given the diverse relationships among map entities, some studies adopt more compound GNNs to enhance graph-based encoders. For example, the HREP model constructs a heterogeneous graph with four types of relations to represent land parcels (Zhou et al., 2023). The MGFN model utilizes human mobility patterns to build a multi-graph, assigning a dedicated GNN to each subgraph (Wu et al., 2022). HRNR (Wu et al., 2020) adopts a hierarchical GNN to capture the connection, structure, and function relations of road segments (Wu et al., 2020). Recent advancements also explore relations across different types of map entities. For instance, HRoad employs hypergraph neural networks to capture the relationships between road segments and land parcels (Zhang and Long, 2023).
@@
Pre-training Tasks. For graph-based models, the corresponding pre-training tasks include three types:
@@
  • Node Feature Inference (NFI). The goal of this task is to infer the features of map entities based on the representations extracted by graph-based encoders. Let $\bm{h}_{i}$ denote the representation of the $i$-th map entity. The NFI task predicts the raw features of the corresponding entity as: $\hat{\bm{f}}_{i}=\phi(\bm{h}_{i})$, where $\hat{\bm{f}}_{i}$ is the predicted feature, and $\phi(\cdot)$ is the prediction function, typically implemented as a multi-layer perceptron (MLP). The learning objective for the NFI task is to minimize the prediction error.
@@
  • Graph Autoencoder (GAu). This task employs an autoencoder to reconstruct the adjacency matrix of relation networks. Given the representation vectors of all map entities, denoted as $\bm{H}$, the GAu task reconstructs the adjacency matrix $\bm{E}$ as $\bm{E}$ as $\hat{\bm{E}}=\phi\left(\bm{H}\bm{H}^{\top}\right)$, where $\hat{\bm{E}}$ is the reconstructed adjacency matrix, and $\phi(\cdot)$ is a neural network acting as the decoder. Here, the graph-based encoder serves as the encoder of the autoencoder. The objective is to minimize the difference between $\hat{\bm{E}}$ and $\bm{E}$.
@@
  • Neighborhood Contrastive Learning (NCL). The intuition behind this task is that connected map entities in a relation network should exhibit similar representations. This task treats connected entities as positive samples and unconnected entities as negative samples. Contrastive learning loss functions are then employed to optimize the encoder parameters.
@@
  • Augmented Graph Contrastive Learning (AGCL). AGCL applies data augmentation-based contrastive learning on map entity relation networks. Given a relation network $G$ and its augmented version $\tilde{G}$, the task treats the representations of the same node across these versions as positive samples, while representations of different nodes are considered negative samples.
@@
In the NFI tasks, most of methods use raw entity features as the supervised labels (Zhang and Long, 2023; Wu et al., 2020; Chen et al., 2024c). In particular, GMEL (Liu et al., 2020) converts the OD flows relation network as in/out-flows features for prediction. In the GAu task, the pre-training method design is closely connected to the structure of GNN in the encoder. For example, HRNR (Wu et al., 2020) uses a three-layer hierarchical GNN, requiring different adjacency matrix reconstruction tasks for each layer. HRoad (Zhang and Long, 2023) adopts a hypergraph in its encoder, so its pre-training step incorporates a hyperedge reconstruction task. In NCL tasks, the primary differences between models lie in the relation networks they utilize. For instance, HREP (Zhou et al., 2023) and ReDCL (Li et al., 2023) connect the geographically nearest entities to incorporate spatial information. In AGCL tasks, the choice of data augmentation is critical. Randomly deleting edges can disrupt essential connections. To mitigate this, some methods use edge weights to determine deletion probabilities (Chang et al., 2023; Zhou et al., 2024).

### v2

# NewREFERENCES

@@ 1. Introduction
In the era of mobile internet, electronic maps have become a foundational platform for a wide range of applications, including intelligent transportation systems and location-based services. An electronic map consists of map entities, such as points of interest (POIs), road segments, and land parcels. These entities encapsulate complex geospatial data, spatial relationships, and geometric topological structures, presenting significant challenges for data representation. The effective representation of map entities has become a critical and enduring research focus within spatiotemporal data analysis and geographic information systems (GIS).
Traditionally, map entities have been stored as structured records in relational databases or geographic file formats, such as PostGIS (Team, 2023), Shapefiles (Environmental Systems Research Institute, 1998), and GeoJSON (Butler et al., 2016). With this representation approach, services relying on map entities are required to develop task-specific models or algorithms to process the entities’ attributes and extract geographic relationships within or between them. However, these task-specific models lack generalizability, limiting their reusability across different applications and scenarios.
In recent years, pre-trained representation learning has emerged as a powerful approach for generating versatile, task-independent data representations. It has achieved remarkable success across various domains, including natural language processing (NLP) (Radford et al., 2018; Peters et al., 2018; Pennington et al., 2014) and computer vision (CV) (Chen et al., 2020; He et al., 2020; Dosovitskiy et al., 2021). The field of spatiotemporal data analysis has also embraced pre-trained representation learning to construct generalized representations of map entities  (Mai et al., 2022; Chen et al., 2024a; Zhang et al., 2024; Li et al., 2024b; Ning et al., 2024), leading to the rise of an emerging research area known as Map Entity Representation Learning (MapRL). However, despite rapid advancements, two major structural challenges require further exploration to unlock the potential of MapRL.
@@
Challenge 2: Lack of Standardized Benchmark. Despite numerous models having been proposed, their performance is often evaluated on different datasets and under varying experimental settings. Unlike the CV and NLP fields, this domain lacks standardized datasets and benchmarks. This absence hinders fair comparisons between models, making it difficult to derive generalizable design principles.
To address the first challenge, we propose a novel taxonomy for MapRL models, namely method-based taxonomy, which organizes the essential components of an MapRL model into four key elements: Map Data, Encoder Models, Pre-training Tasks, and Downstream Tasks. Unlike traditional taxonomies that categorize models by map entity types, method-based taxonomy takes a functional approach, grouping encoder models into three paradigms—Token-based, Graph-based, and Sequence-based—each supporting specialized pre-training tasks. Only downstream tasks depend on the nature of the map entity. A key insight of our taxonomy is that the two core elements—Encoder Models and Pre-training Tasks—are not tightly coupled with specific map entity types. This decoupling of core components from entity-specific constraints fosters a more unified and transferable framework for MapRL, facilitating cross-domain integration and knowledge sharing.
To tackle the second challenge, we introduce the VecCity library, an MapRL model development toolkit guided by our taxonomy. The library organizes map entities and auxiliary data into three atomic file types. Using these atomic files, we format data from nine cities into a unified structure. VecCity further standardizes model implementation with four interface functions covering encoders, pre-training tasks, and downstream tasks. Using these interfaces, we reproduce 21 mainstream MapRL models and conduct evaluations across city datasets, providing key insights for future research.
@@
$\bullet$ Novel Taxonomy. We present a novel taxonomy of MapRL models that goes beyond the traditional classification based on map entities, offering a unified classification system applicable to MapRL models across various types of map entities.
$\bullet$ Easy-to-use Toolkit Library. Guided by our taxonomy, we propose a toolkit, VecCity, which unifies the development process of MapRL models. To the best of our knowledge, VecCity is the first unified model toolkit specifically designed for MapRL.
$\bullet$ Standard Benchmark. VecCity implements 21 mainstream MapRL models and evaluates them on datasets from nine cities, establishing the first unified benchmark for MapRL models and offering valuable insights to guide future research.
@@ 2. Taxonomy of MapRL
In this section, we present the proposed method-based taxonomy. Figure 1 (a) outlines the general framework for MapRL. The training process of a typical MapRL model consists of two key stages: pre-training and fine-tuning. During pre-training, the MapRL model takes map data as input, using an encoder model to transform them into representation vectors. The encoder’s parameters are optimized through pre-training tasks. In the fine-tuning stage, downstream tasks are employed to further refine the encoder. This pipeline defines four core components of an MapRL model: Map Data, Encoder Models, Pre-training Tasks, and Downstream Tasks.
@@ 2.1.1. Map Entities
As illustrated in Fig. 1 (b), an electronic map consists of three geometric elements: Points, Polylines, and Polygons, each representing a distinct type of map entity.
$\bullet$ Points / POI. A point element represents a Point of Interest (POI), identified by a location coordinate, i.e., $(latitude,longitude)$.
$\bullet$ Polylines / Road Segment. A polyline element represents a Road Segment, composed of a sequence of connected line segments, with each endpoint defined by location coordinates.
$\bullet$ Polygons / Land Parcel. A polygon represents a Land Parcel, delineated by a closed sequence of connected line segments.
For each map entity, a set of features describes its properties. For instance, the category feature of a POI differentiates the services it offers. Road segments are characterized by features like speed limits, lane counts, etc. Based on these definitions, we formally define electronic maps and map entities as follows.
Figure 1. Illustration of (a) the framework of a MapRL model and (b) an urban map with different map entities.
@@ 2.1.2. Auxiliary Data
Trajectory data can be divided into two classes based on the type of location samples $l_{tr_{k}}$. $i$) Check-in Trajectories: These use POIs as samples and are collected through social networks or LBS apps. They log a user’s location only upon check-in, resulting in a low sampling rate, often requiring days to capture a single point. MapRL models leverage these trajectories to infer users’ personal preferences. $ii$) Coordinate Trajectories: These rely on geographic coordinates as samples. Collected by ITS through GPS devices or mobile phones, they provide high-frequency sampling along with real-time speed and direction data. These trajectories are aligned with road segments using map matching algorithms (Yang and Gidofalvi, 2018) and facilitate route planning and dynamic state estimation for road segments.
@@
A relation network represents map entities as a directed graph $G=\{V,\bm{E}\}$, where $V=\{v_{1},\ldots,v_{I}\}$ represents the set of vertices corresponding to map entities. The adjacency matrix $\bm{E}\in\mathbb{R}^{I\times I}$ defines the relationships between entities, with each element $e_{ij}$ indicating the relationship from $v_{i}$ to $v_{j}$.
Network relations can be categorized into two types: $i$) Geographical Relations (GR): These leverage spatial relationships between map entities to construct the adjacency matrix. Examples include connections between road segments in a road network (Zhang and Long, 2023) and distances between POIs (Zhao et al., 2017; Du et al., 2019). $ii$) Social Relations (SR): These represent user behaviors, such as mobility trajectories, which are used to construct adjacency matrices. A typical example is Origin-Destination (OD) flows, widely used in MapRL models (Zhang et al., 2020; Zhou et al., 2023), where $e_{ij}$ in the adjacency matrix represents the number of trajectories that start from $v_{i}$ and end at $v_{j}$.
@@ 2.2. Encoder Models and Pre-training Tasks
Encoder models and pre-training tasks are two core components of a pre-trained MapRL model, responsible for converting input map data into representation vectors for map entities. Our taxonomy classifies encoder models into three types: token-based, graph-based, and sequence-based, designed to model features, relation networks, and trajectory sequences of map entities, respectively.
@@ 2.2.1. Token-based Models
Encoders. The token-based encoder generates representation vectors directly from the features of map entities. For discrete features, it applies one-hot encoding, and for continuous features, the value range is divided into consecutive bins, each represented by a one-hot code. Given $K$ features for a map entity, let the one-hot code of the $k$-th feature be $\bm{f}_{k}\in\{0,1\}^{F_{k}}$ (where $F_{k}$ is the dimensionality). The encoder maps each feature to a learnable embedding space using a matrix $\bm{R}_{k}\in\mathbb{R}^{F_{k}\times d}$, generating the entity representation as $\bm{r}=\big{\|}_{k=1}^{K}\bm{f}_{k}^{\top}\cdot\bm{R}_{k}\big{.}$, where $\|_{k=1}^{K}$ denotes the concatenation of embeddings from all features.
Existing token-based methods primarily innovate by refining how they extract meaningful features from map data, which can be classified into three key types: $i$) Spatial Features: These capture the locations and spatial structures of map entities. The most fundamental spatial feature is coordinate information. Hier (Shimizu et al., 2020) enhances spatial representation by embedding hierarchical spatial grids into rasterized maps. $ii$) Temporal Features: The characteristics of a map entity can vary across different times of the day. For instance, a restaurant’s activity level differs between working hours and lunchtime. To model such variations, many token-based encoders divide the day into segments (e.g., hourly or half-hourly) and assign each segment a unique embedding to capture temporal dynamics (Wan et al., 2021; Zhao et al., 2017; Gong et al., 2023; Ma et al., 2024; Liu et al., 2016). $iii$) Semantic Features: These reflect the intrinsic properties of map entities, such as the functional types of POIs or the road types of segments. Some land parcel MapRL methods further enhance semantic features by counting the number of POI categories within a parcel (Zhang et al., 2022).
@@
$\bullet$ Token Relation Inference (TokRI). This task aims to predict the relationship between two map entities based on their representation vectors. Specifically, given the representations $\bm{r}_{i}$ and $\bm{r}_{j}$ of two entities, the TokRI task predicts their relationship as $\hat{y}_{ij}=\phi(\bm{r}_{i},\bm{r}_{j})$, where $\phi(\cdot)$ is the prediction function, and $\hat{y}_{ij}\in[0,1]$ is the predicted probability of a relation existing. The encoder parameters are optimized during pre-training using cross-entropy loss.
@@
$\bullet$ Token Relation Contrastive Learning (TRCL). This task ensures that entities with closer relationships are more likely to have similar representations. Given a representation sample $\bm{c}_{i}$, another sample $\bm{c}_{j}$ is treated as a positive sample if the corresponding entities $v_{i}$ and $v_{j}$ share a close relation; otherwise, $\bm{c}_{j}$ is treated as a negative sample. The encoder parameters are optimized using a contrastive learning loss, such as the InfoNCE loss (Oord et al., 2018).
@@
$\bullet$ Augmented Token Contrastive Learning (AToCL). These tasks use data augmentation techniques to generate different views of a given entity’s representation. A contrastive learning loss is then applied to maximize the similarity between views of the same entity while minimizing the similarity between different samples.
@@
For TokRI and TRCL tasks, the primary innovation in existing models lies in how they construct relations. These methods can be categorized into three classes: $i$) Distance-Based Relations: This approach establishes relations based on spatial proximity. In TokRI tasks, RN2Vec (Wang et al., 2020) defines a relationship between entities if their geographical distance falls below a predefined threshold. In TRCL tasks, Teaser (Zhao et al., 2017) selects the $K$ nearest entities as positive samples for contrastive learning. $ii$) Semantic-Based Relations: Senmant (Huang et al., 2022), Place2V (Yan et al., 2017), and MT-POI (Yao et al., 2023) select entities that have the same or similar semantic features as positive samples for contrastive learning. $iii$) Auxiliary Data-Based Relations: This method leverages additional data, such as trajectories, to define relations. In TokRI tasks, P2Vec (Feng et al., 2017) and Tale (Wan et al., 2021) treat a trajectory as a sentence and the entities within it as words, using the Continuous Bag of Words (CBOW) model (Mikolov et al., 2013) to learn representations for “words” (map entities). In TRCL tasks, CatEM (Bing et al., 2023), R2Vec (Jepsen et al., 2018), and Teaser (Zhao et al., 2017) employ the SkipGram model (Mikolov et al., 2013) with negative sampling to capture contextual relations along trajectories. When trajectory data is unavailable, “virtual trajectories” can be generated via random walks on relation networks (Wang and Li, 2017). Beyond trajectories, ZEMob (Yao et al., 2018) uses OD flows between entities to identify relationships.
@@
For AToCL tasks, the innovation primarily lies in the methods used to generate augmented samples. For instance, ReMVC (Zhang et al., 2022) and ReDCL (Li et al., 2023) enhance contrastive learning by replacing or dropping specific features of the anchor entity to create augmented variants.
@@ 2.2.2. Graph-based Models
Encoders. Graph-based encoders leverage auxiliary data in the form of relation networks (see Definition 3) to capture relationships among map entities. Graph-based encoders typically use the representations generated by token-based encoders as inputs and apply a graph neural network (GNN) to produce refined representation vectors. Specifically, given the map entity set $V=\{v_{1},\ldots,v_{I}\}$, the embeddings generated by the token-based encoder are denoted as $\bm{R}=\{\bm{c}_{1},\ldots,\bm{c}_{I}\}$. The graph-based encoder then takes $\bm{R}$ along with the relation network $G$ to generate the final representations: $\bm{H}=\left\{\bm{h}_{1},\ldots,\bm{h}_{i},\ldots,\bm{h}_{I}\right\}=\mathrm{GNN}\left(\bm{R},G\right),$ where $\bm{h}_{i}$ is the representation vector for the $i$-th map entity, and $\mathrm{GNN}(\cdot)$ is the GNN-based encoder model. Compared to token-based models, graph-based encoders incorporate the relation networks into the entity representations, enriching their expressiveness.
@@
Existing research on graph-based encoders focuses on developing GNNs capable of capturing complex relationships among map entities. Early studies applied standard GNN models, such as graph convolutional networks (Yu et al., 2018; Wu et al., 2019) and graph attention networks (Song et al., 2022; Zhang et al., 2019b), to explore graphical (Chang et al., 2023; Jiang et al., 2023; Jiao et al., 2020) and social relations (Zhang et al., 2020; Liu et al., 2020). Given the diverse relationships among map entities, some studies adopt more compound GNNs to enhance graph-based encoders. For example, the HREP model constructs a heterogeneous graph with four types of relations to represent land parcels (Zhou et al., 2023). The MGFN model utilizes human mobility patterns to build a multi-graph (Wu et al., 2022). HRNR adopts a hierarchical GNN to capture the connection, structure, and function relations of road segments (Wu et al., 2020). Recent advancements also investigate relationships across different types of map entities. For example, HGI (Huang et al., 2023) and HRoad (Zhang and Long, 2023) utilize hypergraphs to capture these inter-entity relationships.
@@
Pre-training Tasks. For graph-based models, the corresponding pre-training tasks include four types:
@@
$\bullet$ Node Feature Inference (NFI). This task aims to predict the features of map entities based on the representations generated by graph-based encoders. Given an entity representation $\bm{h}_{i}$, the NFI task predicts its original features as $\hat{\bm{f}}_{i}=\phi(\bm{h}_{i})$, where $\phi(\cdot)$ is typically an MLP-based prediction function. The objective is to minimize the prediction error.
@@
$\bullet$ Graph Autoencoder (GAu). This task reconstructs the adjacency matrix of relation networks using an autoencoder framework. Given entity representations $\bm{H}$, the adjacency matrix is reconstructed as $\hat{\bm{E}}=\phi(\bm{H}\bm{H}^{\top})$, where $\phi(\cdot)$ acts as a neural decoder. The graph encoder serves as the encoder of the auto-encoder, and the model is optimized to minimize the reconstruction error between $\hat{\bm{E}}$ and the original adjacency matrix $\bm{E}$.
@@
$\bullet$ Neighborhood Contrastive Learning (NCL). The intuition behind this task is that connected map entities in a relation network should exhibit similar representations. This task treats connected entities as positive samples and unconnected entities as negative samples. Contrastive learning loss functions are then employed to optimize the encoder parameters.
@@
$\bullet$ Augmented Graph Contrastive Learning (AGCL). This task introduces data augmentation to contrastive learning in relation networks. Given a relation network $G$ and its augmented variant $\tilde{G}$, the representations of the same node across these versions are treated as positive samples, while different nodes serve as negative samples.
@@
In NFI tasks, most methods leverage raw entity features as supervision signals (Zhang and Long, 2023; Wu et al., 2020; Chen et al., 2024c). In particular, GMEL (Liu et al., 2020) converts the OD flow relation network as labels for prediction. In the GAu task, the design of this task is closely linked to the structure of the GNN encoder. For example, HRNR (Wu et al., 2020) uses a three-layer hierarchical GNN, requiring different adjacency matrix reconstruction tasks per layer. HRoad (Zhang and Long, 2023) adopts a hypergraph in its encoder, so its pre-training step incorporates a hyperedge reconstruction task. In NCL tasks, the primary differences between models lie in the relation networks they utilize. For instance, HREP (Zhou et al., 2023) and ReDCL (Li et al., 2023) connect the nearest entities to incorporate spatial information. In AGCL tasks, the choice of data augmentation is critical. Randomly deleting edges can disrupt essential connections. To mitigate this, some methods use edge weights to determine deletion probabilities (Chang et al., 2023; Zhou et al., 2024).